# Nonadiabatic Origin of Quantum-Metric Effects via Momentum-Space Metric Tensor

### v2

@@ II.3 Effective Mass Corrections for Bound States
where $\bar{G}_{ij}=G_{ij}+\frac{1}{\kappa}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol that vanishes in this example as $G_{ij}$ is constant. $F_{jk}$ corresponds to the artificial magnetic field in the momentum space from Berry curvature. This equation follows the Lagrangian [58]
@@
describing the dynamics of a massive particle under electromagnetic fields on a flat torus with $M=G_{0}+\frac{1}{\kappa}$. By promoting $(p_{i},q^{j})$ to operators with the commutator $\hat{p}_{i},\hat{q}^{j}]=-i\delta_{i}^{j}$, we obtain the quantized Hamiltonian in momentum-space representation. We note that here the canonical momentum $\hat{p}_{i}$ corresponds to the real-space coordinate, ${\bm{x}}=-\frac{1}{\kappa}\dot{\bm{q}}=-\frac{1}{1+\kappa G_{0}}({\bm{p}}+{\bm{A}})$, consistent with the canonical variables in Ref. 30, but with the distinction that it incorporates a nonadiabatic correction from $\kappa G_{0}$.
The topological structure of the Brillouin zone as a torus requires that the $\bm{q}$-space wave functions satisfy periodic boundary conditions. This boundary condition is compatible with the quantization of the Chern number, enabling a direct connection between the electronic states in $\bm{q}$-space to the quantum Hall effects on a torus [84, 85]. The eigenenergies are $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M}=(n+\frac{1}{2})\frac{F\kappa}{1+\kappa G_{0}}$ with each level’s degeneracy equal to $|C|$. The confining potential $\kappa$ corresponds to an effective inverse mass, whereas the nonadiabatic metric contributes to a correction from interband mixing. The interband mixing effect is enhanced by increasing the potential strength that increases $\dot{\bm{q}}$ and thus the interband contributions.
@@
The confining problem can also be generalized to two-body bound states [30] where the effect of $G_{0}$ can be enhanced as $\kappa$ originates from the Coulomb interaction strength. Here we consider two wave packets in flat bands. We assume that the wave packets are well separated, and they interact through a softened Coulomb interaction in real space. The two-particle Lagrangian is
@@
If the two particles are different, like an electron and a hole, the wave function of the relative coordinate is only restricted by the periodic boundary condition. If they are indistinguishable, the two-body wavefunctions need to be antisymmetric for fermions or symmetric for bosons. This leads to another boundary condition in addition to the periodic boundary conditions. Therefore, while the degeneracy of eigen-energy levels obtained by solving equation (18) is the Chern number $|C|$ ($C\neq 0$), only those states satisfying both boundary conditions are allowed solutions.
A natural realization of the two-body problem is the excitons formed by electrons and holes in flat bands [86, 30, 87]. The potential relationship with Cooper pairs is interesting and warrants further investigation [88, 89, 90, 91].
II.4 Metric Tensor and Forced Geodesic Equation
The nonadiabatic metric also underscores the importance of the $\bm{q}$-space picture, instead of the real space, where $\bm{q}$ is taken as a generalized coordinate to describe the wave packet dynamics. This enables us to study the emergent gravity and emergent mass in momentum space, which will be discussed in this and following sections, separately.
$G_{ij}(\bm{q})$ in the Lagrangian defines a metric tensor in the tangent vector space of the generalized velocity $\dot{\bm{q}}$ defining a line element $ds^{2}=G_{ij}dq^{i}dq^{j}$. In contrast to the quantum metric, which defines the quantum distance $ds_{Q}$ between two wave functions $|u_{\bm{q}}\rangle$ and $|u_{\bm{q}+d\bm{q}}\rangle$ at $\bm{q}$ and $\bm{q}+d\bm{q}$ via $ds_{Q}^{2}=1-|\langle u_{\bm{q}}|u_{\bm{q}+d\bm{q}}\rangle|^{2}$ [3], the nonadiabatic metric defines the geometric distance between the momentum-space centers of two wave packets. A $\bm{q}$-dependent $G_{ij}$ signifies a curved manifold. Conversely, in cases where $G_{ij}$ is constant across the momentum space, it corresponds to an effective mass in a flat $\bm{q}$-space, simplifying the dynamics to a Euclidean framework.
where $\Gamma^{l}_{jk}$ is the Christoffel symbol of the second kind
$\Gamma^{i}_{jk}=\frac{1}{2}G^{il}\left(\frac{\partial G_{lj}}{\partial q^{k}}+\frac{\partial G_{kl}}{\partial q^{j}}-\frac{\partial G_{jk}}{\partial q^{l}}\right),$ (20)
$G^{ji}G_{ik}=\delta^{j}_{k}$ with $\delta^{j}_{k}$ the Kronecker delta. When the right-hand side of equation (19) vanishes, the equation reduces to the geodesic equation for a free particle navigating a curved manifold with metric tensor $G_{ij}$. A nonzero right-hand side indicates a particle subject to forces from three sources (see Fig. 3): a scalar potential from the band dispersion $E$, a gauge field from the Berry curvature $F_{ij}$, and an external force from the wave packet velocity $\dot{x}_{j}$. This equation enables the study of gravity in condensed matter setups [92]. It is noted that equation (19) differs from the recent study on momentum space gravity [93, 94, 95] where $G^{ij}$ and $\ddot{\bm{q}}$ are absent and $\Gamma_{i,jk}$ instead of $\Gamma^{i}_{jk}$ is used.
the equations of motion find their $\bm{q}$-space representation:
where $\bar{G}_{ij}=G_{ij}+\Xi_{ij}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol. With information of $\dot{\bm{q}}$, $x_{i}$ can be solved using algebraic equations. The metric tensor becomes dynamical with $\dot{\bar{G}}_{ij}=\frac{\partial G_{ij}}{\partial q_{k}}\dot{q}_{k}+\frac{\partial\Xi_{ij}}{\partial\dot{q}_{k}}\ddot{q}_{k}$.
III Discussions
@@
We introduce a geometric framework that extends the study of Bloch electron dynamics in a scalar potential beyond adiabatic Berry-phase effects. We show that the leading-order corrections to the adiabatic dynamics, the nonadiabatic effects, are elegantly characterized by the nonadiabatic metric—a momentum-space metric tensor in semiclassical wave packet dynamics, distinguishing them from the conventional quantum metric. This metric unifies the nonlinear and nonadiabatic responses by introducing geometric and geodesic velocities beyond the anomalous velocity from the Berry curvature, enabling further generalization. It also modifies the quantum electron dynamics, manifested as a metric in momentum space. A constant nonadiabatic metric influences the flat-band electron dynamics in a harmonic potential by introducing a correction to the effective mass for the generalized coordinate $\bm{q}$. Depending on the origin of the harmonic potential, this mass correction influences the electronic properties in static potentials, Wigner crystal [78, 81], or flat-band excitons [30]. In general, the nonadiabatic metric is not a constant, which endows momentum space with a curved geometry, making the electronic dynamics a forced geodesic motion in momentum space. This enables the study of momentum-space gravity.
IV Methods
IV.1 Integrate out $c_{n}$
Given the Lagrangian
$L$ $=\sum_{n}c_{n}^{*}i\dot{c}_{n}-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}+\Lambda_{[00]}+\sum_{n\neq 0}-c_{n}^{*}c_{n}\left[\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]})\right]+(c_{0}^{*}c_{n}\Lambda_{[0n]}+c.c.)$
where $c_{0}$ is the reference-band amplitude and $c_{n}$ are dynamical variables similar to $\bm{x}_{c}$ and $\bm{q}_{c}$. The normalization implies $c_{0}=1+\mathcal{O}(c_{n\neq 0}^{2})$ and the phase of $c_{0}$ contributes a total derivative and can be gauged away. For simplicity, we define
$\Delta_{n}\equiv\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]}),\qquad J_{n}\equiv c_{0}\Lambda_{[n0]},~n\neq 0.$ (23)
The action $S=\int dtL$ is quadratic in the excited-band amplitudes $c_{n\neq 0}$. Varying S with respect to $c_{n}^{*}$ gives the saddle-point equation
$(i\partial_{t}-\Delta_{n})c_{n}=-J_{n}\qquad n\neq 0.$ (24)
[…]
$S_{\rm eff}[\bm{q}_{c},\bm{x}_{c}]$ $=\int dtL_{0}(\bm{q}_{c},\bm{x}_{c})$
where $L_{0}(\bm{q}_{c},\bm{x}_{c})\equiv-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c,0}+\Lambda_{[00]}$.
When parameters $\bm{q}_{c}$ vary slowly, we expand the inverse operator:
$G_{n}$ $=(i\partial_{t}-\Delta_{n})^{-1}\delta(t-t^{\prime})$
$=\left[-\frac{1}{\Delta_{n}}+\frac{i}{\Delta_{n}^{2}}\partial_{t}+\mathcal{O}(\partial_{t}^{2})\right]\delta(t-t^{\prime}),$ (26)
and keep terms up to second order in time derivatives. Using $J_{n}=c_{0}\Lambda^{n0}$ and integrating by parts, equation (25) gives a local effective Lagrangian
$\mathcal{L}_{\rm eff}=L_{0}+\sum_{n\neq 0}\frac{J_{n}^{*}J_{n}}{\Delta_{n}}+\mathcal{O}(\partial_{t}).$ (27)
By substituting $J_{n}=c_{0}\dot{\bm{q}}_{c}\cdot\bm{A}_{[n0]}$ into the above equation and neglecting the Berry phase corrections to the band gap, i.e., $|\Lambda_{[nn]}-\Lambda_{[00]}|\ll|\omega_{[n0]}|$ and $\Delta_{n}\simeq\omega_{[n0]}$, we obtain the effective Lagrangian as shown in equation (6).
V Acknowledgments

### v3

@@ IV.2 Massive Particle Dynamics in Flat Bands
where $\bar{G}_{ij}=G_{ij}+\frac{1}{\kappa}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol that vanishes in this example as $G_{ij}$ is constant. The group velocity vanishes as $E$ is a constant and $F_{jk}$ corresponds to the artificial magnetic field in the momentum space from Berry curvature. This equation follows the Lagrangian [52]
@@
describing the dynamics of a massive particle under electromagnetic fields on a flat torus with $M=G_{0}+\frac{1}{\kappa}$. By promoting $(p_{i},q^{j})$ to operators with the commutator $\hat{p}_{i},\hat{q}^{j}]=-i\delta_{i}^{j}$, we obtain the quantized Hamiltonian in momentum-space representation. We note that here the canonical momentum $\hat{p}_{i}$ corresponds to the real-space coordinate, ${\bm{x}}=-\frac{1}{\kappa}\dot{\bm{q}}=-\frac{1}{1+\kappa G_{0}}({\bm{p}}+{\bm{A}})$, consistent with the canonical variables in Ref. 76, but with the distinction that it incorporates a nonadiabatic correction from $\kappa G_{0}$.
The topological structure of the Brillouin zone as a torus requires that the $\bm{q}$-space wave functions satisfy periodic boundary conditions. This boundary condition is compatible with the quantization of the Chern number, enabling a direct connection between the electronic states in $\bm{q}$-space to the quantum Hall effects on a torus [84, 85]. The eigenenergies are $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M}=(n+\frac{1}{2})\frac{F}{1/\kappa+G_{0}}$ with each level’s degeneracy equal to $|C|$. The confining potential $\kappa$ corresponds to an effective inverse mass that is modified by the nonadiabatic metric $G_{0}$ from interband mixing. The effect of $G_{0}$ is enhanced by increasing $\kappa$, i.e., increasing the potential strength, which can increase $\dot{\bm{q}}$ and thus the interband contributions.
@@
IV.3 Two-Body Bound States in Flat Bands
The confining problem can also be generalized to two-body bound states [76] where the effect of $G_{0}$ can be enhanced as $\kappa$ originates from the Coulomb interaction strength. Here we consider two wave packets in flat bands. We assume that the two particles live in two layers well separated by a distance $x_{0}$, and they interact through a Coulomb interaction in real space. The two-particle Lagrangian is
@@ IV.3 Two-Body Bound States in Flat Bands
If the two particles are different, like an electron and a hole, the wave function of the relative coordinate is only restricted by the periodic boundary condition. Here we consider ideal situations to illustrate the effects of the nonadiabatic metric. Their effects in more realistic conditions deserve detailed study in the future, which are naturally related to excitons formed by electrons and holes in flat bands [86, 76, 87]. The potential relationship with Cooper pairs is interesting and warrants further investigation [88, 89, 90, 91].
V Summary
@@
We introduce a geometric framework that extends the study of Bloch electron dynamics in a scalar potential beyond adiabatic Berry-phase effects. We show that the leading-order nonadiabatic effects are elegantly characterized by the nonadiabatic metric—a momentum-space metric tensor that is related to but distinct from the quantum metric. This nonadiabatic metric unifies the nonlinear and nonadiabatic responses by introducing geometric and geodesic velocities beyond the anomalous velocity from the Berry curvature. The geometric velocity is related to the metric tensor, whereas the geodesic velocity is related to the Christoffel symbol. The nonadiabatic metric also modifies the quantum dynamics of electrons, manifested as a metric in momentum space. A constant nonadiabatic metric influences the flat-band electron dynamics in a harmonic potential by introducing a correction to the effective mass for the generalized coordinate $\bm{q}$. In general, the nonadiabatic metric is not a constant, which endows momentum space with a curved geometry, making the electronic dynamics a forced geodesic motion in momentum space.
Acknowledgments